OpenStudy (anonymous):

What is an equationof the line that passes through the points (-6,-2) and (4,-2)

OpenStudy (anonymous):

First find the slope \[ m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] once you have that use point - slope form \[y-y _{1}=m(x-x _{1})\] and turn it into an equation of a line, being y = mx+b.

OpenStudy (amriju):

y=-2

hero (hero):

Plot both points on an xy coordinate plane then draw a line through them.